# Can We Infer Confidential Properties of Training Data from LLMs?

### v3

@@ Abstract
Large language models (LLMs) are increasingly fine-tuned on domain-specific datasets to support applications in fields such as healthcare, finance, and law. These fine-tuning datasets often have sensitive and confidential dataset-level properties — such as patient demographics or disease prevalence—that are not intended to be revealed. While prior work has studied property inference attacks on discriminative models (e.g., image classification models) and generative models (e.g., GANs for image data), it remains unclear if such attacks transfer to LLMs. In this work, we introduce PropInfer, a benchmark task for evaluating property inference in LLMs under two fine-tuning paradigms: question-answering and chat-completion. Built on the ChatDoctor dataset, our benchmark includes a range of property types and task configurations. We further propose two tailored attacks: a prompt-based generation attack and a shadow-model attack leveraging word frequency signals. Empirical evaluations across multiple pretrained LLMs show the success of our attacks, revealing a previously unrecognized vulnerability in LLMs.
@@ 2 Related Work
Property inference: Property Inference Attack (PIA) was first described by ateniese2013hackingsmartmachinessmarter , as follows: given two candidate training data distributions $\mathcal{D}_{1},\mathcal{D}_{2}$ and a target model, the adversary tries to guess which training distribution (out of $\mathcal{D}_{1},\mathcal{D}_{2}$) is the target model trained on. Typically, the two candidate distributions only differ in the marginal distribution of a binary variable, such as gender ratio. A major portion of past work on property inference focuses on discriminative models ateniese2013hackingsmartmachinessmarter ; ganju2018property ; chase2021propertyinferencepoisoning ; suri2024dissectingdistributioninference ; zhang2021leakagedatasetpropertiesmultiparty ; 10136150 ; here the attacks mainly rely on training meta-classifiers on some representations to predict target ratio. For example, in the white-box setting, ateniese2013hackingsmartmachinessmarter ; ganju2018property use model weights as the input of the meta-classifier to predict the correct distribution. In the grey-box setting, where the adversary have access to the training process and some auxiliary data, suri2022formalizingestimatingdistributioninference ; suri2024dissectingdistributioninference use model outputs such as loss or probability vector as inputs to the meta-classifier.
@@ 3.1 Large Language Models Fine-Tuning
A large language model (LLM) predicts the likelihood of a sequence of tokens. Given input tokens $t_{0},...,t_{i-1}$, the language model parameterized by parameters $\theta$, $f_{\theta}$, outputs the distribution of the possible next token $f_{\theta}(t_{i}|t_{0},...,t_{i-1})$. Pre-training LLMs on large-scale corpora enables them to develop general language understanding and encode broad world knowledge. In pre-training, the LLM is trained to maximize the likelihood of unlabeled text sequences. Each training sample is a document comprising a sequence of tokens, and the objective is to minimize the negative log-likelihood: $\mathcal{L}(\theta)=-\log\sum_{i=1}^{k}f_{\theta}(t_{i}|t_{0},...t_{i-1})$ where $k$ is the total number of tokens in the document.
@@
    Supervised Fine-Tuning (SFT; (radford2018improving, ; ouyang2022training, )). SFT minimizes the negative log-likelihood of the output tokens conditioned on the instruction and input. This approach focuses on learning the mapping from $(I,x)$ to $y$ and is commonly used in question-answering tasks. The objective is: $\mathcal{L}_{\rm SFT}(\theta)=-\log\sum_{i=1}^{l}f_{\theta}(y_{i}|I,x,y_{0},...,y_{i-1}).$
@@
    Causal Language-Modeling Fine-Tuning (CLM-FT;  (radford2018improving, )). Different from SFT, CLM-FT follows the pre-training paradigm and minimizes the loss over all tokens in the concatenated sequence $t$ of instruction, input, and output $(I,x,y)$. This method treats the full sequence autoregressively, making it suitable for tasks involving auto-completion for both user and chatbot. The objective is $\mathcal{L}_{\rm CLM-FT}(\theta)=-\log\sum_{i=1}^{k}f_{\theta}(t_{i}|t_{0},...t_{i-1}).$
@@ 5.2 Shadow-Model Attack with Word Frequency under Grey-Box Setting
Constructing new shadow attacks with word frequency. The choice of the shadow feature functions $F$ plays an important role in the success of the attack. While previous work relies on loss or probability vector suri2022formalizingestimatingdistributioninference ; suri2024dissectingdistributioninference , some studies have shown that these features may not be the most effective way to measure the performance of the LLMs carlini2021extractingtrainingdatalarge ; duan2024membershipinferenceattackswork . Hence, we propose a new feature specific to the LLM setting, i.e. word frequency. Our attack is based on the intuition that certain properties may strongly correlate with the appearance of specific words in the text. As a result, models fine-tuned on datasets with different property distributions may exhibit distinct word patterns in their generations.
@@ 6.2 Results
Gender property inference. Table 1 presents the results of our attacks on the gender property inference task for models fine-tuned in both Q&A Mode and Chat-Completion Mode. We highlight two main observations. First, in Q&A Mode, our word-frequency attack significantly outperforms both baselines and our BB generation attack. Second, in Chat-Completion Mode, the BB generation attack achieves the best performance, with the word-frequency attack performing closely behind – both substantially outperforming the baselines.
The strong performance of the word-frequency attack, particularly in Q&A Mode, can be attributed to two factors. First, it operates under a stronger threat model by leveraging an auxiliary dataset, unlike the black-box methods. Second, word frequency provides a more effective signal than the perplexity-based baseline. In Appendix A.3, we include examples of the keyword list used in our word-frequency attack, which reveals interpretable correlations with the gender property.
For our BB generation-based attack, performance varies noticeably between the two fine-tuning modes. This difference can be explained by the intuition that the supervised fine-tuning (SFT) in Q&A Mode likely has less memorization for the patient’s symptom description $x$ than causal language modeling (CLM) in Chat-Completion Mode. Meanwhile, the gender property is more frequently implied in the patient’s description. Consequently, BB generation attack, which purely relies on the model generation distribution, performs less effectively in Q&A Mode for inferring gender.
Medical diagnosis property inference. Table 2 presents the results of our attacks on the medical diagnosis property inference task for models fine-tuned in both Q&A Mode and Chat-Completion Mode. We highlight two main observations that are consistent across both fine-tuning modes and all three LLMs: First, our BB generation attack achieves strong performance and consistently outperforms both baselines across all three diagnosis attributes. Second, the attack performs relatively worse on the childbirth attribute compared to mental disorder and digestive disorder.
The relatively lower performance on the childbirth attribute may be explained by the results of the Generation w/o FT baseline. We suspect this is due to the cultural sensitivity of childbirth-related topics (e.g., pregnancy, abortion), which may have led to safety training during pretraining that suppresses the generation of such content. As a result, the pretrained model’s output distribution is likely the most misaligned with the fine-tuned target distribution for this property, reflected by the highest MAE among the three attributes. This might limits the effectiveness of our attack.
@@
6.3 Ablation study
We conduct an ablation study to assess the impact of key hyperparameters in both of our proposed attacks. presents the results of this study for the gender attribute using the LLaMA-3 model. Ablation results for additional model architectures and properties can be found in Appendix A.3.
For the word-frequency attack, we examine two factors when testing with the Q&A Mode: the number of selected keywords $d$ and the total number of shadow models $k_{1}\cdot k_{2}$. As shown in Figure 2(a), the optimal number of keywords lies between 30 and 35. Using too few keywords may result in weak signals, while too many can introduce noise and overwhelm the meta attack model, given a limited number of shadow models. Figure 2(b) shows that increasing the number of shadow models improves performance, as it provides more training data for the meta model, enhancing its generalization.
@@ 7 Discussion
Possible Defenses. Even though it is impossible to provide confidentiality for all properties of a dataset and still produce an useful model, in most practical cases only a small subset of properties are confidential, and these are often largely unrelated to the intended use of the model. For example, the income-level of the customers is unrelated to answering customer service questions.
One plausible defense strategy is to subsample the training data, resulting in a dataset more closely aligned with a known public prior. Although subsampling can mitigate property inference attacks at their source, it may also compromise model utility by limiting the amount of effective training data. An alternative approach is to reweight the training data, either by duplicating certain samples or adjusting their weights in the loss. This method preserves exposure to the full dataset while implicitly altering the learned distribution. However, its effectiveness as a defense remains to be validated.
@@ 8 Conclusion
Limitation and future work. Firstly, although our attack has a high success rate in inferring the proportion of mental disorder and digestive disorder, it has a low success rate in childbirth; therefore, a natural future work is to propose better attacks to investigate whether there are privacy leakages for childbirth. Secondly, while subsampling can mitigate property inference at its source, it is not ideal when the dataset is limited or the training task requires large amount of data. Hence, more future works on better defenses are needed to protect data confidentiality.
@@ Appendix A Technical Appendices and Supplementary Material
  • In Section A.2, we present the details of the shadow model attack with perplexity.
@@
  • In Section A.4, we present the experimental setup, including model performance, fine-tuning parameters and details of the baseline attacks.
@@ A.1 Labeling process:
A.2 Shadow-model attack with perplexity.
@@ A.3 Ablation Study
  • For the word-frequency attack, we study the effect of the number of keywords $d$ and the number of shadow models $k_{1}\cdot k_{2}$ on the attack performance.
@@
Figure 3 shows the ablation study in the Q&A mode; the optimal number of keywords for word frequency attack varies between different architectures. For the Llama1 model, the optimal number of keywords is less than 5; for example, when $d=3$, the chosen keywords are "spotting", "female" and "scanty", where "spotting" and "female" are gender-indicated words. For the Llama3 model, the optimal number of keywords lies between 30 and 35; when $d=30$, some examples of the chosen keywords are "cigarette", "smoked", "nifedipine", "gynecomastia", "epigastric", where "gynecomastia" is gender-indicated word and "cigarette" and "smoked" are more common in male than in female. For the pythia model, the optimal number of keywords lies between 65 and 75; when $d=65$, some examples of the chosen keywords are "pelvic", "vaginal, "indigestion", "painkiller", "urinary" and "backache", where "vaginal" is gender-indicated word. We observe that the chosen keywords as well as the number of keywords are very distinct between models; we suspect the reason is that the pre-training data distribution and the model architecture is different for three base models, hence it may have an effect of the generated text distributions.
@@
Figure 4 shows the ablation study in the Chat-Completion mode. For Llama1 model, the optimal number of keywords is between 363-6; when $d=5$, the chosen keywords are "his", "her", "he", "female", and "she", where all chosen keywords are clearly gender-indicated. For the Llama3 model, the optimal keywords are between 353-5; when $d=5$, the chosen keywords are "penile, "female", "scrotal", "masturbating" and "erection", where all chosen keywords are gender-indicated. For the Pythia model, the mean absolute error is less than $5\%$ for $d<70$, which shows that the attack performance is effective; when $d=5$, the chosen keywords are "scrotum", "penis", "foreskin", "glans" and "female". We observe that in the Chat-Completion mode, all the selected keywords are clearly gender-indicated and with a very small number of keywords, the word-frequency based shadow model attack achieves an effective performance.
@@
A.4 Experiment Setup
Experiment compute resources:
Target Model performance
As shown in table 4 and 5, we evaluate the performance of the target models using the BERT score[30], following [17]. In particular, we choose a subset with size $500$ from a separate test dataset, iCliniq dataset, provided by [17]. We generate outputs given the inputs using greedy decoding and calculate the BERT score between the generated texts and the labels. We observe that the fine-tuned Pythia model, as well as the Pythia base model, sometimes outputs an empty string, hence we only calculate the BERT score between non-empty outputs and its corresponding labels. The performance of these models is similar to the performance reported in the paper [17].
Dataset Model Precision Recall F1 Score
Table 4: Target model evaluation using BERT score in Q&A mode.
Dataset Model Precision Recall F1 Score
We query each target model with three different prompts as follows.
[…]
  • Prompt 1 = Instruction + "Hi, Chatdoctor, I have a medical question."
  • Prompt 2 = Instruction + "Hi, doctor, I have a medical question."
  • Prompt 3 = Instruction + "Hi Chatdoctor, here is my question."
Baseline attack: shadow-model attack with perplexity.
Details of this attack is included in A.2
Baseline direct asking prompts:
we use the following prompts to query the target models:
[…]

### v4

@@ Abstract
Large language models (LLMs) are increasingly fine-tuned on domain-specific datasets to support applications in fields such as healthcare, finance, and law. These fine-tuning datasets often have sensitive and confidential dataset-level properties — such as patient demographics or disease prevalence—that are not intended to be revealed. While prior work has studied property inference attacks on discriminative models (e.g., image classification models) and generative models (e.g., GANs for image data), it remains unclear if such attacks transfer to LLMs. In this work, we introduce PropInfer, a benchmark task for evaluating property inference in LLMs under two fine-tuning paradigms: question-answering and chat-completion. Built on the ChatDoctor dataset, our benchmark includes a range of property types and task configurations. We further propose two tailored attacks: a prompt-based generation attack and a shadow-model attack leveraging word frequency signals. Empirical evaluations across multiple pretrained LLMs show the success of our attacks, revealing a previously unrecognized vulnerability in LLMs. We release our code at github.com/PengrunH/Property_inference_attack_LLM.
@@ 2 Related Work
Property inference. Property Inference Attack (PIA) was first described by ateniese2013hackingsmartmachinessmarter , as follows: given two candidate training data distributions $\mathcal{D}_{1},\mathcal{D}_{2}$ and a target model, the adversary tries to guess which training distribution (out of $\mathcal{D}_{1},\mathcal{D}_{2}$) is the target model trained on. Typically, the two candidate distributions only differ in the marginal distribution of a binary variable, such as gender ratio. A major portion of past work on property inference focuses on discriminative models ateniese2013hackingsmartmachinessmarter ; ganju2018property ; chase2021propertyinferencepoisoning ; suri2024dissectingdistributioninference ; zhang2021leakagedatasetpropertiesmultiparty ; 10136150 ; here the attacks mainly rely on training meta-classifiers on some representations to predict target ratio. For example, in the white-box setting, ateniese2013hackingsmartmachinessmarter ; ganju2018property use model weights as the input of the meta-classifier to predict the correct distribution. In the grey-box setting, where the adversary have access to the training process and some auxiliary data, suri2022formalizingestimatingdistributioninference ; suri2024dissectingdistributioninference use model outputs such as loss or probability vector as inputs to the meta-classifier.
@@ 3.1 Large Language Models Fine-Tuning
A large language model (LLM) predicts the likelihood of a sequence of tokens. Given input tokens $t_{0},...,t_{i-1}$, the language model parameterized by parameters $\theta$, $f_{\theta}$, outputs the distribution of the possible next token $f_{\theta}(t_{i}|t_{0},...,t_{i-1})$. Pre-training LLMs on large-scale corpora enables them to develop general language understanding and encode broad world knowledge. In pre-training, the LLM is trained to maximize the likelihood of unlabeled text sequences. Each training sample is a document comprising a sequence of tokens, and the objective is to minimize the negative log-likelihood: $\mathcal{L}(\theta)=-\sum_{i=1}^{k}\log f_{\theta}(t_{i}|t_{0},...t_{i-1})$ where $k$ is the total number of tokens in the document.
@@
    Supervised Fine-Tuning (SFT; (radford2018improving, ; ouyang2022training, )). SFT minimizes the negative log-likelihood of the output tokens conditioned on the instruction and input. This approach focuses on learning the mapping from $(I,x)$ to $y$ and is commonly used in question-answering tasks. The objective is: $\mathcal{L}_{\rm SFT}(\theta)=-\sum_{i=1}^{l}\log f_{\theta}(y_{i}|I,x,y_{0},...,y_{i-1}).$
@@
    Causal Language-Modeling Fine-Tuning (CLM-FT;  (radford2018improving, )). Different from SFT, CLM-FT follows the pre-training paradigm and minimizes the loss over all tokens in the concatenated sequence $t$ of instruction, input, and output $(I,x,y)$. This method treats the full sequence autoregressively, making it suitable for tasks involving auto-completion for both user and chatbot. The objective is $\mathcal{L}_{\rm CLM-FT}(\theta)=-\sum_{i=1}^{k}\log f_{\theta}(t_{i}|t_{0},...t_{i-1}).$
@@ 5.2 Shadow-Model Attack with Word Frequency under Grey-Box Setting
Constructing new shadow attacks with word frequency. The choice of the shadow feature functions $F$ plays an important role in the success of the attack. While previous work relies on loss or probability vector suri2022formalizingestimatingdistributioninference ; suri2024dissectingdistributioninference , some studies have shown that these features may not be the most effective way to measure the performance of the LLMs carlini2021extractingtrainingdatalarge ; duan2024membershipinferenceattackswork . We alternatively focus on another feature specific to the LLM setting, word frequency, which has been parallelly studied for the literature of membership inference attack kaneko2024samplingbasedpseudolikelihoodmembershipinference ; meeus2025canarysechoauditingprivacy . Our attack is based on the intuition that certain properties may strongly correlate with the appearance of specific words in the text. As a result, models fine-tuned on datasets with different property distributions may exhibit distinct word patterns in their generations.
@@ 6.2 Main Results
Gender property inference.
Table 1 presents the results of our attacks on the gender property inference task for models fine-tuned in both Q&A Mode and Chat-Completion Mode. We highlight two main observations: First, in Q&A Mode, our word-frequency attack significantly outperforms both baselines and our BB generation attack. Second, in Chat-Completion Mode, the BB generation attack achieves the best performance, with the word-frequency attack performing closely behind – both substantially outperforming the baselines.
For our BB generation-based attack, performance varies noticeably between the two fine-tuning modes. For example, for the Llama-1 model, the Q&A mode yields a high MAE of $17.5\%$, whereas the CC mode achieves a much lower MAE of $2.55\%$. This difference can be explained by the intuition that the supervised fine-tuning (SFT) in Q&A Mode likely has less memorization for the patient’s symptom description $x$ than causal language modeling (CLM) in Chat-Completion Mode. Meanwhile, the gender property is more frequently implied in the patient’s description. Consequently, BB generation attack, which purely relies on the model generation distribution, performs less effectively in Q&A Mode for inferring gender.
Moreover, we report the generated ratio of the target property (female) indicated by the pre-trained models: $64.2\%$ (Llama-1), $67.7\%$ (Pythia), $62.6\%$ (Llama-3). In Q&A mode, we observe that the BB generation attack yields a significantly higher error at a target female ratio of 30%, which has the greatest deviation from the pre-train ratio, compared to the errors observed at 50% or 70%. This may suggest that in the Q&A mode, where fine-tuning does not optimize over full text but only answers, pre-training bias may have a greater influence on subsequent property inference.
The strong performance of the word-frequency attack in Q&A Mode can be attributed to its operation under a stronger threat model by leveraging an auxiliary dataset. Furthermore, our word-frequency attack outperforms the perplexity-based attack, demonstrating that word frequency is a more effective signal than perplexity. Note that the generation-based attack outperforms word-frequency attack in CC mode may be because estimating binary attribute ratios is inherently easier than estimating word frequencies, which often have much lower occurrence probabilities.
Medical diagnosis property inference.
Table 2 presents the results of our attacks on the medical diagnosis property inference task for models fine-tuned in both Q&A Mode and Chat-Completion Mode. We highlight two main observations that are consistent across both fine-tuning modes and all three LLMs: First, our BB generation attack achieves strong performance and consistently outperforms both baselines across all three diagnosis attributes. Second, the attack performs relatively worse on the childbirth attribute compared to mental disorder and digestive disorder.
The relatively lower performance on the childbirth attribute may be explained by the results of the Generation w/o FT baseline. For example, the exact ratios of three medical properties from the pre-trained Llama-3 model are $1.65\%$ (mental disorder), $8.03\%$ (digestive disorder) and $0.279\%$ (childbirth). We observe that the pre-train ratio for childbirth is notably lower than the pre-train ratio for the other two properties. We suspect this is due to the cultural sensitivity of childbirth-related topics (e.g., pregnancy, abortion), which may have led to safety training during pretraining that suppresses the generation of such content. As a result, the pretrained model’s output distribution is likely the most misaligned with the fine-tuned target distribution for this property, reflected by the highest MAE among the three attributes. This might limits the effectiveness of our attack.
@@
6.3 Additional studies
Empirical analysis of selected keywords.
To better understand the word-frequency attack, we provide examples of the keyword list chosen by the shadow models in Table 3. For each word, we compute the correlation coefficient between its occurrence frequency and the corresponding female ratio. A correlation coefficient close to 0 indicates no correlation, while values near 1 or -1 indicate strong positive or negative correlation. We find that many selected keywords exhibit strong correlations with the target female ratio, supporting the effectiveness of word frequency as a signal. Notably, the correlations are generally stronger when the target model is in CC mode compared to QA mode—aligning with the performance differences observed in Table 1, where word-frequency attacks perform better under CC mode than QA mode. More details are provided in Appendix A.3.
Table 3: Top 5 chosen Keywords and the corresponding Correlation Coefficients for Llama-1 Models
Model Keywords (Correlation Coefficients)
Llama-1 (CC mode) his (0.956-0.956), himself (0.934-0.934), her ($0.947$), he (0.950-0.950), female ($0.960$)
Llama-1 (QA mode) female ($0.785$), reluctant ($0.774$), spotting ($0.787$), scanty ($0.821$), ovaries ($0.782$)
Ablation study of our attacks.
We conduct an ablation study to assess the impact of key hyperparameters in both of our proposed attacks. We include the results of this study for the gender attribute using the LLaMA-3 model. Additional ablation studies can be found in Appendix A.3.
@@ 6.3 Additional studies
For the word-frequency attack, we examine two factors when testing with the Q&A Mode: the number of selected keywords $d$ and the total number of shadow models $k_{1}\cdot k_{2}$. As shown in Figure 2(a), the optimal number of keywords lies between 30 and 35. Using too few keywords may result in weak signals, while too many can introduce noise and overwhelm the meta attack model, given a limited number of shadow models. Figure 2(b) shows that increasing the number of shadow models improves performance, as it provides more training data for the meta-model, enhancing its generalization.
Additionally, we conduct ablation studies on relaxation on some assumption of the shadow models. To test the assumption of access to the same pre-trained model, we evaluate a setting where the adversary uses Llama-3 for training shadow models but targets a Llama-1 model. MAE results for both Q& A and CC modes are shown in table 4. We observe that the knowledge of the pre-trained model has small effects in CC mode, but notable effects in the QA mode. This might be because with the CC mode, the model is learned to fully mimic the texts in the fine-tuning dataset, disregarding what the pre-trained model is. However, in QA mode, while the model learns to answer questions based on the fine-tuning data, the generation of the questions themselves may still rely heavily on the pre-trained model, leading to greater divergence in the resulting question–answer pairs. More experiments can be found in Appendix  A.3.
@@
Table 4: Testing shadow model’s assumption on pre-trained model architecture
Q&A Mode Chat-Completion Mode
30 50 70 30 50 70
Knowing the pre-trained model $6.86\%$ $12.96\%$ $6.82\%$ $3.44\%$ $0.27\%$ $6.26\%$
Without knowing the model $20.23\%$ $0.24\%$ $20.00\%$ $3.85\%$ $0.17\%$ $15.31\%$
@@ 7 Discussion
Possible defenses. Even though it is impossible to provide confidentiality for all properties of a dataset and still produce a useful model, in most practical cases only a small subset of properties are confidential, and these are often largely unrelated to the intended use of the model. For example, the income-level of the customers is unrelated to answering customer service questions.
One plausible defense strategy is to subsample the training data, resulting in a dataset more closely aligned with a known public prior. Although subsampling can mitigate property inference attacks at their source, it may also compromise model utility by limiting the amount of effective training data. An alternative approach is to adjust the output sampling methods: since the attack depends on model generations, we implement a simple defense that adjust the temperature parameter in the final softmax layer, so that the generation does not reflect the ground-truth distribution; the experiment details are put in Appendix A.4. However, such a defense may break down if the adversary is aware of the default temperature setting or has access to the model weights. A deeper investigation into the limitations of this defense, as well as the development of more robust mitigation strategies, is an important direction for future work.
8 Acknowledge
This work was supported in part by the ONR under grants N000142412304 and N00014-25-1-2116, by the NSF under grants CIF-2402817 and CNS-2241100, and by the ARO-MURI under grant W911NF2110317.
We thank Yiwei Fu for assistance with checking the GPT-based labeling used in this work.
@@ 9 Conclusion
Limitation and future work. Firstly, although our attack has a high success rate in inferring the proportion of mental disorder and digestive disorder, it has a low success rate in childbirth; therefore, a natural future work is to propose better attacks to investigate whether there are privacy leakages for childbirth. Secondly, datasets from other domains can also be relevant under the property inference threat model — for example, capturing the distribution of opinions in a news dataset. Extending the benchmark to include such datasets represents an important direction for strengthening the evaluation. Lastly, while subsampling can mitigate property inference at its source, it is not ideal when the dataset is limited or the training task requires large amount of data. Hence, more future works on better defenses are needed to protect data confidentiality.
@@ Appendix A Technical Appendices and Supplementary Material
  • In Section A.2, we present the experimental setup, including fine-tuning details, model performance, and details of the baseline attacks.
@@
  • In Section A.4, we include a simple defense mechanisms.
@@ A.1 Labeling process:
A.2 Experiment Setup
Experiment compute resources:
Target Model performance
As shown in table 5 and 6, we evaluate the performance of the target models using the BERT score[32], following [18]. In particular, we choose a subset with size $500$ from a separate test dataset, iCliniq dataset, provided by [18]. We generate outputs given the inputs using greedy decoding and calculate the BERT score between the generated texts and the labels. We observe that the fine-tuned Pythia model, as well as the Pythia base model, sometimes outputs an empty string, hence we only calculate the BERT score between non-empty outputs and its corresponding labels. The performance of these models is similar to the performance reported in the paper [18].
Dataset Model Precision Recall F1 Score
Table 5: Target model evaluation using BERT score in Q&A mode.
Dataset Model Precision Recall F1 Score
We query each target model with three different prompts as follows.
[…]
  • Prompt 1 = Instruction + "Hi, Chatdoctor, I have a medical question."
  • Prompt 2 = Instruction + "Hi, doctor, I have a medical question."
  • Prompt 3 = Instruction + "Hi Chatdoctor, here is my question."
Baseline attack: shadow-model attack with perplexity.
@@ A.2 Experiment Setup
Baseline direct asking prompts:
we use the following prompts to query the target models:
[…]
@@ A.3 Ablation Study
  • For the word-frequency attack, we study the effect of the number of keywords $d$ and the number of shadow models $k_{1}\cdot k_{2}$ on the attack performance. We also report the selected keyword list for all target models.
@@
Figure 3 shows the ablation study in the Q&A mode; the optimal number of keywords for word frequency attack varies between different architectures. For the Llama1 model, the optimal number of keywords is less than 5; for example, when $d=3$, the chosen keywords are "spotting", "female" and "scanty", where "spotting" and "female" are gender-indicated words. For the Llama3 model, the optimal number of keywords lies between 30 and 35. For the pythia model, the optimal number of keywords lies between 65 and 75. We observe that the chosen keywords as well as the number of keywords are very distinct between models; we suspect the reason is that the pre-training data distribution and the model architecture is different for three base models, hence it may have an effect of the generated text distributions.
@@
Figure 4 shows the ablation study in the Chat-Completion mode. For Llama1 model, the optimal number of keywords is between 363-6; when $d=5$, the chosen keywords are "his", "her", "he", "female", and "she", where all chosen keywords are strongly gender-indicated. For the Llama3 model, the optimal keywords are between 353-5; when $d=5$, the chosen keywords are "penile, "female", "scrotal", "masturbating" and "erection", where all chosen keywords are gender-indicated. For the Pythia model, the mean absolute error is less than $5\%$ for $d<70$, which shows that the attack performance is effective; when $d=5$, the chosen keywords are "scrotum", "penis", "foreskin", "glans" and "female". We observe that in the Chat-Completion mode, all the selected keywords are strongly gender-indicated and with a very small number of keywords, the word-frequency based shadow model attack achieves an effective performance.
@@
Additionally, We put the keyword list chosen by the shadow models in Table 7. For each word, we compute the correlation coefficient between its occurrence frequency and the corresponding female ratio. A correlation coefficient close to 0 indicates no correlation, while values near 1 or -1 indicate strong positive or negative correlation, respectively. Notably, for all models, we find that the correlations are generally stronger when the target model is in CC mode compared to QA mode—aligning with the performance differences observed in Table 1, where word-frequency attacks perform better under CC mode than QA mode.
Table 7: Top Keywords and Correlation Coefficients for Different Models and Modes
Model (Mode) Keywords (Correlation Coefficients)
Llama-1 (CC) his (0.956-0.956), himself (0.934-0.934), her ($0.947$), he (0.950-0.950), female ($0.960$),
him (0.933-0.933), prostate (0.931-0.931), she ($0.954$), son (0.929-0.929), daughter ($0.932$)
Pythia (CC) scrotum (0.941-0.941), he (0.916-0.916), penis (0.944-0.944), foreskin (0.934-0.934),
male (0.928-0.928), glans (0.940-0.940), female ($0.968$), masturbate (0.919-0.919),
masturbation (0.921-0.921), tip (0.917-0.917)
Llama-3 (CC) penile (0.940-0.940), erect (0.935-0.935), penis (0.939-0.939), scrotum (0.935-0.935),
female ($0.980$), scrotal (0.941-0.941), males (0.935-0.935), masturbating (0.945-0.945),
erection (0.945-0.945), erectile (0.934-0.934)
Llama-1 (QA) female ($0.785$), reluctant ($0.774$), spotting ($0.787$), scanty ($0.821$), ovaries ($0.782$)
Pythia (QA) pelvic ($0.738$), recurring ($0.730$), football (0.740-0.740), bland ($0.760$),
indigestion ($0.740$), bothering ($0.822$), uti ($0.746$), point ($0.743$),
presenting (0.751-0.751), smear ($0.753$)
Llama-3 (QA) nifedipine (0.786-0.786), readings (0.788-0.788), yielding (0.765-0.765), squats (0.788-0.788),
analogs (0.791-0.791), smoke (0.773-0.773), particular ($0.847$), cigarette (0.774-0.774),
quit (0.810-0.810), regularly (0.832-0.832)
@@
A.4 Defenses
we implemented a simple defense: since the attack depends on model generations, we adjusted the temperature parameter $T$ in the final softmax layer. Note that $T>1$ makes the model’s output more balanced among all tokens and $T<1$ makes the model’s output more concentrated on the high-probability tokens.
We empirically evaluate this defense on the BB-generation attack using LLaMA-3 fine-tuned in CC mode. The table below reports the average predicted ratios inferred by the attack under different temperature settings, given fine-tuning datasets with varying female ratios.
Table 9: Effect of Different $T$ Values on Target Female Ratios
$T=1$ (no defense) $T=0.5$ $T=1.5$ $T=3.5$
Target Female Ratio = 0.3 0.355 0.200 0.415 0.4585
Target Female Ratio = 0.5 0.516 0.447 0.542 0.5657
Target Female Ratio = 0.7 0.6958 0.8686 0.6453 0.6866
We observe that without defense ($T=1$ ), the attack predict the target female ratio mostly correctly. When the temperature is altered ($T\neq 1$) , the predicted ratios have larger error, indicating that adjusting decoding settings can serve as a simple yet effective defense against black-box sampling-based attacks. However, such a defense may break down if the adversary is aware of the default temperature setting or has access to the model weights. A deeper investigation into the limitations of this defense, as well as the development of more robust mitigation strategies, is an important direction for future work.